OpenStudy (anonymous):

hi! i have a problem... i don't have enough experience with geometry but hopefully someone will help.

OpenStudy (mrnood):

what is the question?

OpenStudy (anonymous):

@MrNood can u help me with my question pls?

OpenStudy (anonymous):

Quadrilateral ABCD is located at A(−2, 2), B(−2, 4), C(2, 4), and D(2, 2). The quadrilateral is then transformed using the rule (x + 7, y − 1) to form the image A'B'C'D'. What are the new coordinates of A', B', C', and D'?

OpenStudy (mrnood):

it tells you that a value of x is transformed to x + 7 y is transformed to y-1 the first point has x= -2 so the transrmend point is x= (-2+7) = 5 the first point has y = 2 so the transformed point is y= (2-1) =1 so A' = (5,1) do the same for the other 3 points

OpenStudy (anonymous):

ok thank you so much!! :~)